OpenStudy (anonymous):

solve for x: 4^(3x+8)=12 please show steps

OpenStudy (anonymous):

step one take the log \[(3x+8)\ln(4)=\ln(12)\] step two distribute \[3\ln(4)x+8\ln(4)=\ln(12)\] subtract \[3\ln(4)x=\ln(12)-8\ln(4)\] divide \[x=\frac{\ln(12)-8\ln(4)}{3\ln(4)}\]

OpenStudy (anonymous):

loved your work in "it takes a thief"!